# On completeness of Foliated structures, and Null Killing fields

### v1

@@ 7.3.1. Case of $\dim G\geq 3$
It appears from Proposition 7.3 that this case corresponds to a null Killing field $V$ having a dense orbit in $M$ (in which case, all orbits of $V$ are dense). We saw in Proposition 7.3 that in this case, $M$ is a flat $3$-torus, which is then in particular complete, and $\dim G=3$.
@@ 7.3.2. Case of $\dim G=1$
In this paragraph, we assume that $\dim G=1$, i.e. the flow of $V$ is periodic. Then $M$ is a Seifert fibration. For simplicity, we will assume that the $\mathbb{S}^{1}$-action of $V$ is free. Then $V$ is tangent to the fibres of a fibre bundle $\pi:M\to\Sigma$, where $\Sigma$ is diffeomorphic to a $2$-torus.
@@
The following lemma is a converse for Lemma 7.5 above.
@@
Let $\pi:M\to\Sigma$ be an $\mathbb{S}^{1}$-bundle over a $2$-torus $\Sigma$, given by an $\mathbb{S}^{1}$ free action. Denote by $V$ the vector field on $M$ generated by the $\mathbb{S}^{1}$-action (it is tangent to the fibers). Let $(X^{*},Y^{*})$ be two vector fields on $\Sigma$ such that $[X^{*},Y^{*}]=f^{*}Y^{*}+h^{*}X^{*}$, with $f^{*},h^{*}\in C^{\infty}(\Sigma,\mathbb{R})$. There exist vector fields $(X,Y)$ on $M$ together with a Lorentzian metric $g$, such that
@@
  • $[X,Y]=fY+hX+\mu V$, with $f,h,\mu\in C^{\infty}(M,\mathbb{R})$ are $V$-invariant functions such that $f,h$ project to $f^{*},h^{*}$ respectively.
@@
With the above reduction, the equation on $b(t)$ for an isotropic geodesic reads
@@
(1) $\Gamma(t)$ does not cross any band. Indeed, if $\gamma$ crosses a band of $X$, then there is a point where $\gamma$ is tangent to $X$. But this means that $\alpha=0$. On the other hand, if $\gamma$ crosses a band of $Y$, then there is a point where $\gamma$ is tangent to $Y$. This means that $b(t_{0})=0$ for some $t_{0}\in I$. By Observation 7.13, this yields $b$ identically zero. In both cases, this contradicts our assumptions on $\gamma$.
@@
Step 1: We first prove that $\Gamma$ remains in the half band delimited by $l$ and $l_{1}$, so that $x(t)$ is bounded on $\Gamma$. The geodesic can leave the half band either by crossing the band, or by making a turnaround. Looking at the frame field $(X^{*},Y^{*})$ on the boundary leaves (see Figure 1), it appears that the first behaviour cannot occur. More precisely, if $\Gamma$ crosses the half band, then there is a point $t_{0}$ such that $x(t_{0})=\pi$ and $x^{\prime}(t_{0})\geq 0$. But Equation (3) yields $b(t_{0})\leq 0$, which is impossible. On the other hand, it follows from Lemma 7.14 that $\Gamma$ does not make turnarounds. The claim follows.
@@
Comment 7.16.
The general completeness question in dimension $3$ in the case of a periodic action reduces to Question 7.12. It would be interesting to investigate this question. It seems to us that the completeness depends on the dynamics of the vector fields $X^{*}$ and $Y^{*}$ defined on the torus.
@@ 7.3.3. Case of $\dim G=2$
Fact 7.17.
Let $(M,V)$ be a compact Lorentzian $3$-manifold with an equicontinuous null Killing field $V$, such that the closure of the flow of $V$ in ${\sf{Isom}}(M)$ is isomorphic to $\mathbb{T}^{2}$. Assume that the $\mathbb{T}^{2}$-action is free. Then
@@
    If all the $\mathbb{T}^{2}$-leaves are degenerate, then the manifold is complete.
@@
    If all the $\mathbb{T}^{2}$-leaves are Lorentzian, then the manifold is complete.
@@
(1) Let $K$ be a Killing field transversal to $V$, hence necessarily spacelike. Let $Y$ be the (unique) null vector field such that $g(Y,K)=0$ and $g(Y,V)=1$. Then $(V,K,Y)$ is a $V$-invariant frame field. Let $\gamma$ be a geodesic, and write $\dot{\gamma}(t)=a(t)V+b(t)K+c(t)Y$. We have that $c(t)=g(\dot{\gamma}(t),V)$ and $g(\dot{\gamma}(t),K)=b(t)g(K,K)$ are constant, and that $a(t)$ satisfies a differential equation $\dot{a}(t)=F(\gamma(t))$, where $F(p):=\alpha\frac{\nabla_{Y_{p}}g(K,K)}{g(K(p),K(p))^{2}}$, $\alpha\in\mathbb{R}$, is a bounded function on $M$ (see the note after Example 7.20, where we give the explicit computations leading to the equation on $a(t)$). Since $K$ is spacelike and $M$ is compact, the component $b(t)$ is a bounded function. Moreover, $a(t)$ is bounded on finite time intervals. Thus, the components of $\dot{\gamma}$ are bounded on finite time intervals, and the completeness follows.
(2) To prove completeness, we will show that there is a timelike Killing field generated by the action of $\mathbb{T}^{2}$. This action is free and proper, so the quotient map $\pi:M\to M/\mathbb{T}^{2}\simeq\mathbb{S}^{1}$ is a $\mathbb{T}^{2}$-principal bundle. Denote by $\mathcal{F}$ the $2$-dimensional foliation tangent to the fibers, and by $x$ the coordinate on the base space $\mathbb{S}^{1}$. Since the $\mathbb{T}^{2}$-action is isometric, it preserves the unitary vector field $T$ tangent to the orthogonal distribution $T\mathcal{F}^{\perp}$. It follows that the $\mathbb{T}^{2}$-action commutes with the flow of $T$. Denote by $C_{0}$ (resp. $C_{x}$) the nullcone field of the Lorentzian metric on $F_{0}$ (resp. on $F_{x}$). Let $x\in\mathbb{S}^{1}$, the (non oriented) angle $\alpha_{x}$ of the timecone delimited by $C_{x}$ is well-defined; indeed, since $\mathbb{T}^{2}$ is abelian, the Lorentzian metric on the fiber preserved by the $\mathbb{T}^{2}$-action is flat, so the nullcone field $C_{x}$ is parallel along the fiber. Moreover, since the parameter $x$ varies in a compact set, and the nullcone field varies continuously, $\alpha_{x}$ has a minimum $>0$. Now, to guarantee the existence of a timelike Killing field, we proceed as follows. For two leaves $F_{0}$ and $F_{x}$, $x\in\mathbb{S}^{1}$, there is a diffeomorphism $\phi^{t_{x}}_{T}:F_{0}\to F_{x}$, where $\phi^{t}_{T}$ is the flow of $T$. The pullback of the nullcone field on $F_{x}$ by the flow of $T$ is another nullcone field on $F_{0}$, whose (pullbacked) angle is $\alpha_{x}$. Since $M$ has a null Killing field $V$ commuting with $T$, the nullcone field $C_{0}$ and the pullbacks of the $C_{x}$, for all $x\in\mathbb{S}^{1}$, have a common (invariant) null vector field $V_{|F_{0}}$. This means that we can choose a timelike vector ${\mathfrak{z}}$ tangent to $F_{0}$ at some point $p\in F_{0}$, with the property that all its push-forwards by the $T$-flow are also timelike. Since the (isometric) action of $\mathbb{T}^{2}$ is transitive on the fibers, there exists a Killing field $K$ such that $K(p)={\mathfrak{z}}$. Since $K$ is $T$-invariant and all $d\phi_{T}^{t}({\mathfrak{z}})$ are timelike, $K$ is timelike along the orbit of $T$ through $x$ (which intersects all the fibers). Finally, the $\mathbb{T}^{2}$-invariance of $K$ proves that it is timelike everywhere. ∎
@@
In the proof of Item $(2)$ above, the flow of $T$ does not necessarily preserve the null cone fields of the leaves. In fact, those are preserved by the flow of $T$ if and only if the flow induces a conformal diffeomorphism between the leaves. The foliation $\mathcal{F}^{\perp}$ is totally geodesic (normal-(((which is equivalent to $\mathcal{F}$ being transversally Riemannian). And $\mathcal{F}$ is umbilic (which is equivalent to $\mathcal{F}^{\perp}$ being transversally conformal, see [24, Fact 2.1])normal-))), and since $\dim\mathcal{F}^{\perp}=1$, this is equivalent to the flow of $T$ preserving the conformal structure of the leaves of $\mathcal{F}$. When it is the case, by [24, Proposition 2.4], this leads to the warped product structure in Example 7.19 below.
@@
Consider $\mathbb{R}^{3}$ with the Lorentzian metric $g_{f}=f(z)(dx^{2}-dy^{2})+dz^{2}$, where $f(z)>0$ and $l$-periodic. When $f$ is non-constant, $(\mathbb{R}^{3},g_{f})$ is non-flat. Moreover, the isometry group is, up to some discrete factor, $\mathbb{R}^{2}\times l\mathbb{Z}$. Hence, the quotient $M_{f}:=\mathbb{R}^{3}/\Gamma$ by a uniform lattice $\Gamma\subset\mathbb{R}^{3}$ is a compact Lorentzian manifold with connected component of the isometry group isomorphic to $\mathbb{T}^{2}$. The isometry group contains the flow of a null Killing field, and one can choose $\Gamma$ in such a way that this flow is dense in $\mathbb{T}^{2}$.
@@
Consider $\mathbb{R}^{3}$ with the Lorentzian metric $g=(dudv+\cos(t)du^{2})+dt^{2}$. Let $V:=\partial_{v},U:=\partial_{u},T:=\partial_{t}$. Both $U$ and $V$ are Killing fields (with $V$ null) generating an $\mathbb{R}^{2}$-isometric action with Lorentzian flat leaves. The null cone field is spanned by $U$ and $U-\frac{1}{2}cos(t)V$. Since $[T,U-\frac{1}{2}\cos(t)V]=\frac{1}{2}\sin(t)V$ (normal-(((here, $[T,V]=0,[T,U]=0$)normal-))), the line field $\mathbb{R}(U-cos(t)V)$ is not preserved by $T$. So the flow of $T$ does not preserve the null cone field, and the decomposition $(\mathbb{R}^{2}\times\mathbb{R},(dudv+\cos(t)du^{2})+dt^{2})$ is therefore not a warped product. Note that this example is in fact a pp-wave, written in Brinkmann coordinates.
Note: Explicit computations leading to the geodesic equation in Fact 7.17
Let $(V,K,Y)$ be a $V$-invariant frame field such that $V$ and $K$ are Killing, and the scalar products $g(K,V),g(K,Y),g(V,V),g(Y,Y)$ and $g(Y,V)$ are constant. We claim that all Christoffel symbols $\Gamma_{**}^{Y}$ are null but $\Gamma_{KK}^{Y}=\frac{1}{2}\nabla_{Y}g(K,K)$. The only non-straightforward computation is $\Gamma_{KV}^{Y}$. Using that $K$ and $V$ are both Killing, that the scalar products $g(K,Y),g(Y,V)$ are constant, and that $[V,K]=0$, we compute the scalar products:
$\bullet\;\;g([K,Y],K)=g(\nabla_{K}Y,K)-g(\nabla_{Y}K,K)$ $=-g(\nabla_{K}K,Y)-g(\nabla_{Y}K,K)$
$=g(\nabla_{Y}K,K)-g(\nabla_{Y}K,K)=0$
$\bullet\;\;g([K,Y],Y)=g(\nabla_{K}Y,Y)-g(\nabla_{Y}K,Y)=0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;$
$\bullet\;\;g([K,Y],V)=g(\nabla_{K}Y,V)-g(\nabla_{Y}K,V)$ $=-g(\nabla_{K}V,Y)+g(\nabla_{V}K,Y)$
$=g(\nabla_{V}K,Y)+g(\nabla_{V}K,Y)=0.$
This proves that $[K,Y]=0$. Using this, compute $g(\nabla_{Y}V,K)=-g(\nabla_{K}V,Y)=g(\nabla_{K}Y,V)=g(\nabla_{Y}K,V)=-g(\nabla_{Y}V,K)$, which proves that $g(\nabla_{Y}V,K)=0$. From this one infers that $\Gamma_{KV}^{Y}=\Gamma_{VK}^{Y}=0$. On the other hand, it is clear that $\Gamma_{KK}^{Y}=\frac{1}{2}\nabla_{Y}g(K,K)$. Finally, the vanishing of the other Christoffel symbols follow straightforwardly from the constancy of the scalar products in the lemma.

### v2

@@ 7.3.1. Case of $\dim G\geq 3$
By Proposition 7.1, this case corresponds to a null Killing field $V$ having a dense orbit in $M$ (in which case, all orbits of $V$ are dense). As shown in Proposition 7.1, $M$ is, in this case, a flat $3$-torus, which is therefore complete, with $\dim G=3$.
@@ 7.3.2. Case of $\dim G=1$
In this paragraph, we assume that $\dim G=1$, i.e. the flow of $V$ is periodic. Then $M$ is a Seifert fibration. For simplicity, we will assume that the $\mathbb{S}^{1}$-action of $V$ is free. Then $V$ is tangent to the fibers of a fiber bundle $\pi:M\to\Sigma$, where $\Sigma$ is diffeomorphic to a $2$-torus.
@@
The following lemma is a converse to Lemma 7.7 above.
@@
Let $\pi:M\to\Sigma$ be an $\mathbb{S}^{1}$-bundle over a $2$-torus $\Sigma$, given by an $\mathbb{S}^{1}$ free action. Denote by $V$ the vector field on $M$ generated by the $\mathbb{S}^{1}$-action (it is tangent to the fibers). Let $(X^{*},Y^{*})$ be two vector fields on $\Sigma$ such that $[X^{*},Y^{*}]=f^{*}Y^{*}+h^{*}X^{*}$, with $f^{*},h^{*}\in C^{\infty}(\Sigma,\mathbb{R})$. There exist vector fields $X,Y$ on $M$ together with a Lorentzian metric $g$ on $M$, such that
@@
  • $[X,Y]=fY+hX+\mu V$, where $f,h,\mu\in C^{\infty}(M,\mathbb{R})$ are $V$-invariant functions on $M$, such that $f$ and $h$ project to $f^{*}$ and $h^{*}$ respectively.
@@
With the above reduction, the equation on $b(t)$ for a null geodesic reads
@@
(1) $\Gamma(t)$ does not cross any band. Indeed, if $\gamma$ crosses a band of $X$, then there is a point where $\gamma$ is tangent to $X$. But this means that $\alpha=0$. On the other hand, if $\gamma$ crosses a band of $Y$, then there is a point where $\gamma$ is tangent to $Y$. This means that $b(t_{0})=0$ for some $t_{0}\in I$. By Observation 7.15, this yields $b$ identically zero. In both cases, this contradicts our assumptions on $\gamma$.
@@
Step 1: We first prove that $\Gamma$ remains in the half band delimited by $l$ and $l_{1}$, so that $x(t)$ is bounded on $\Gamma$. The geodesic can leave the half band either by crossing the band, or by making a turnaround. Looking at the frame field $(X^{*},Y^{*})$ on the boundary leaves (see Figure 1), it appears that the first behavior cannot occur. More precisely, if $\Gamma$ crosses the half band, then there is a point $t_{0}$ such that $x(t_{0})=\pi$ and $x^{\prime}(t_{0})\geq 0$. But Equation (4) yields $b(t_{0})\leq 0$, which is impossible. On the other hand, it follows from Lemma 7.16 that $\Gamma$ does not make turnarounds. The claim follows.
@@
Comment 7.18.
The general completeness question in dimension $3$ in the case of a periodic action reduces to Question 7.14. This question deserves further investigation. It seems to us that the completeness depends on the dynamics of the vector fields $X^{*}$ and $Y^{*}$ defined on the torus (we hope to come back to it in a future work).
@@ 7.3.3. Case of $\dim G=2$
Proposition 7.19.
Let $(M,V)$ be a compact Lorentzian $3$-manifold with an equicontinuous null Killing field $V$, such that the closure of the flow of $V$ in ${\sf{Isom}}(M)$ is isomorphic to $\mathbb{T}^{2}$. Assume that the $\mathbb{T}^{2}$-action is free, and let $\mathcal{F}$ be the $2$-dimensional foliation defined by this action. Then
@@
    If all the $\mathcal{F}$-leaves are degenerate, then the manifold is complete.
@@
    If all the $\mathcal{F}$-leaves are Lorentzian, then the manifold is complete.
@@
(1) The group $\mathbb{T}^{2}$ provides an abelian subalgebra of Killing fields that generates the totally geodesic lightlike foliation tangent to $V^{\perp}$. By [13, Theorem 3], we conclude that $M$ is a (compact) plane wave and, therefore, complete.
(2) To prove completeness, we will show that there is a timelike Killing field generated by the action of $\mathbb{T}^{2}$. This $\mathbb{T}^{2}$-action is free and proper, so the quotient map $\pi:M\to M/\mathbb{T}^{2}\simeq\mathbb{S}^{1}$ is a $\mathbb{T}^{2}$-principal bundle, whose fibers are given by the leaves of $\mathcal{F}$. Denote by $x$ the coordinate on the base space $\mathbb{S}^{1}$. Since the $\mathbb{T}^{2}$-action is isometric, it preserves the unitary vector field $T$ tangent to the orthogonal distribution $T\mathcal{F}^{\perp}$. It follows that the $\mathbb{T}^{2}$-action commutes with the flow of $T$. Denote by $C_{0}$ (resp. $C_{x}$) the nullcone field of the Lorentzian metric on $F_{0}$ (resp. on $F_{x}$). Let $x\in\mathbb{S}^{1}$, the (non oriented) angle $\alpha_{x}$ of the timecone delimited by $C_{x}$ is well-defined; indeed, since $\mathbb{T}^{2}$ is abelian, the Lorentzian metric on the fiber preserved by the $\mathbb{T}^{2}$-action is flat, so the nullcone field $C_{x}$ is parallel along the fiber. Moreover, since the parameter $x$ varies in a compact set, and the nullcone field varies continuously, $\alpha_{x}$ has a minimum $>0$. Now, to guarantee the existence of a timelike Killing field, we proceed as follows. For two leaves $F_{0}$ and $F_{x}$, $x\in\mathbb{S}^{1}$, there is a diffeomorphism $\phi^{t_{x}}_{T}:F_{0}\to F_{x}$, where $\phi^{t}_{T}$ is the flow of $T$. The pullback of the nullcone field on $F_{x}$ by the flow of $T$ is another nullcone field on $F_{0}$, whose (pullbacked) angle is $\alpha_{x}$. Since $M$ has a null Killing field $V$ commuting with $T$, the nullcone field $C_{0}$ and the pullbacks of the $C_{x}$, for all $x\in\mathbb{S}^{1}$, have a common (invariant) null vector field $V_{|F_{0}}$. This means that we can choose a timelike vector ${\mathfrak{z}}$ tangent to $F_{0}$ at some point $p\in F_{0}$, with the property that all its push-forwards by the $T$-flow are also timelike. Since the (isometric) action of $\mathbb{T}^{2}$ is transitive on the fibers, there exists a Killing field $K$ such that $K(p)={\mathfrak{z}}$. Since $K$ is $T$-invariant and all $d\phi_{T}^{t}({\mathfrak{z}})$ are timelike, $K$ is timelike along the orbit of $T$ through $x$ (which intersects all the fibers). Finally, the $\mathbb{T}^{2}$-invariance of $K$ proves that it is timelike everywhere. ∎
@@
In the proof of Item $(2)$ of Proposition 7.19, the flow of $T$ does not necessarily preserve the null cone fields of the $\mathcal{F}$-leaves. In fact, the latter are preserved by the flow of $T$ if and only if the flow induces a conformal diffeomorphism between the leaves. The foliation $\mathcal{F}^{\perp}$ is totally geodesic ((((equivalently, $\mathcal{F}$ is transversely Riemannian)))). And since $\dim\mathcal{F}^{\perp}=1$, $\mathcal{F}^{\perp}$ is transversely conformal if and and only if the flow of $T$ preserves the conformal structure of the leaves of $\mathcal{F}$. In this case, $\mathcal{F}$ is umbilic by [35, Fact 2.1]. When this occurs, by [35, Proposition 2.4], this leads to the warped product structure in Example 7.21 below.
We mentioned in the proof of Proposition 7.19 that the metrics in Item (1) are plane waves. Let us now look more closely to the situation in Item (2), and see which kind of metrics we can have there.
@@
Consider $\mathbb{R}^{3}$ with the Lorentzian metric $g_{f}=f(z)(dx^{2}-dy^{2})+dz^{2}$, where $f>0$ and $\tau$-periodic. When $f$ is non-constant, $(\mathbb{R}^{3},g_{f})$ is non-flat. The connected component of the isometry group of $g_{f}$ contains ${\mathsf{Sol}}$, which acts by preserving the Lorentzian flat planes spanned by $\partial_{x}$ and $\partial_{y}$. Its maximal normal abelian subgroup acts by translations in the $x$ and $y$ coordinates. And it contains a one parameter subgroup acting by $A^{t}:(x,y,z)\mapsto(e^{t}x,e^{-t}y,z)$. Since $f$ is $\tau$-periodic, there is in addition an isometric action of $\mathbb{Z}$ by $(x,y,z)\mapsto(x,y,z+n\tau),n\in\mathbb{Z}$. Clearly, this action commutes with that of ${\mathsf{Sol}}$. Define a quotient $M_{f}:=\mathbb{R}^{3}/\Gamma$ by a discrete subgroup $\Gamma\subset{\mathsf{Sol}}\times\mathbb{Z}$ of the form $\mathbb{Z}^{2}\times\mathbb{Z}$, where $\mathbb{Z}^{2}$ is a lattice of the maximal normal abelian subgroup of ${\mathsf{Sol}}$. Then:
1) The isometry group of $M$ is non-compact. Indeed, there is $\lambda>0$ for which $A^{\lambda}$ preserves the lattice $\Gamma$, hence induces a non-trivial action on the quotient. This leads to an isometric action of $\mathbb{Z}$ on $M$, coming from $A^{\lambda}$.
2) By [8, Lemma 2.1], for a generic $f$, the connected component of the isometry group of $g_{f}$ is exactly ${\mathsf{Sol}}$. In this case, $M$ is a compact Lorentzian manifold with connected component of the isometry group isomorphic to $\mathbb{T}^{2}$. The isometry group contains the flow of a null Killing field, and one can choose $\Gamma$ in such a way that this flow is dense in $\mathbb{T}^{2}$, leading to the situation in Proposition 7.19, Item (2).
Proposition 7.22.
Consider $X=\mathbb{R}^{3}$ equipped with a warped product metric $g_{f}=f(z)dxdy+dz^{2}$, where $f$ is periodic. Then $(X,g_{f})$ admits a parallel null vector field $V$, i.e. the warped product is a Brinkmann manifold, if and only if $g_{f}$ is flat.
Proof.
First observe that ${\sf{Isom}}^{\mathsf{o}}(X)$ contains ${\mathsf{Sol}}$, which acts by preserving the flat Lorentzian leaves. Assume $X$ is a Brinkmann manifold. Let $V$ be the parallel null vector field of the Brinkmann structure. Since the warped product is non-flat, the function $f$ is non-constant. Therefore, it follows from [30, Chapter 7, Proposition 35] that the Killing fields generated by one-parameter groups of ${\mathsf{Sol}}$ cannot be parallel. Hence, the one-parameter group generated by $V$ in ${\sf{Isom}}^{\mathsf{o}}(X)$ is not contained in ${\mathsf{Sol}}$. So the dimension of ${\sf{Isom}}^{\mathsf{o}}(X)$ is at least $4$. By [8, Theorem 4.1], $X$ must be locally homogeneous. On the other hand, since $f$ is periodic, $X$ admits a compact quotient $M=X/\Gamma$ with a non-compact isometry group (see Example 7.21). This compact quotient is locally homogeneous. By [8, Theorem C point (2)], it is flat, and then $g_{f}$ is also flat. ∎
As noted in Remark 7.20, the decomposition coming from an isometric action of $\mathbb{T}^{2}$ with Lorentzian leaves is a warped product decomposition if and only if $\mathcal{F}^{\perp}$ is transversely conformal. In the next example, however, $\mathcal{F}^{\perp}$ is not transversely conformal. Moreover, the metric is not isometric to a metric $g_{f}$ as in Example 7.21.
@@
Consider $X:=\mathbb{R}^{3}$ with the Lorentzian metric $g=(dudv+\cos(t)du^{2})+dt^{2}$. Let $V:=\partial_{v},U:=\partial_{u},T:=\partial_{t}$. Both $U$ and $V$ are Killing fields (with $V$ null) that generate an $\mathbb{R}^{2}$-isometric action with Lorentzian flat leaves. The null cone field is spanned by $U$ and $U-\frac{1}{2}cos(t)V$. Since $[T,U-\frac{1}{2}\cos(t)V]=\frac{1}{2}\sin(t)V$, the line field $\mathbb{R}(U-cos(t)V)$ is not preserved by $T$. So the flow of $T$ does not preserve the null cone field, and the decomposition $(\mathbb{R}^{2}\times\mathbb{R},(dudv+\cos(t)du^{2})+dt^{2})$ is therefore not a warped product with respect to our decomposition. Observe that this example is a (non-flat) pp-wave, written in Brinkmann coordinates. So, by Proposition 7.22, it is not isometric to a warped product of the form $g_{f}$, independently of the decomposition. Moreover, $\mathbb{R}^{2}$ acts isometrically by translations in the $u$ and $v$ coordinates. Thus, taking the quotient of $X$ by a discrete group of isometries generated by a lattice in $\mathbb{R}^{2}$ along with the isometry $\gamma:(u,v,t)\mapsto(u,v,t+2\pi)$ results in a compact quotient where the connected component of the isometry group is precisely $\mathbb{T}^{2}$, with Lorentzian $\mathcal{F}$-leaves.
The authors do not know at this stage if there are examples in Item (2) where the universal cover is neither a warped product nor a Brinkmann spacetime. A broader question is: what compact Lorentzian $3$-manifolds arise when we assume the existence of an equicontinuous null Killing field?